OpenStudy (itsbribro):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2?

OpenStudy (sweetburger):

do you know the point slope equation?

OpenStudy (itsbribro):

no

OpenStudy (itsbribro):

iv'e learnt about it but i have bad memory

OpenStudy (sweetburger):

(y-y1) = m(x-x1) is the point slope equation you literally just plug in the numbers given in the question

OpenStudy (sweetburger):

so you take your y coordinate and put in in for y1 you put your x coordinate in the x1 and the slope value goes in the m slot

OpenStudy (sweetburger):

so it would like this (y-5) = 2(x-3)

OpenStudy (itsbribro):

ok so it would look like this?\[(y-5)=2(x-3)\]

OpenStudy (sweetburger):

yes there you go

OpenStudy (sweetburger):

gj!

OpenStudy (itsbribro):

thanks

OpenStudy (sweetburger):

np

OpenStudy (chosenmatt):

yup